# Secure numerical simulations using fully homomorphic encryption

### v2

@@ 4.3.3 Rewriting the finite difference methods in 2D
Note that $\bm{u}^{n+1}$ is a matrix now, with coefficients $u_{ij}$ at each node location. Consequently, we also need to use the matrix representation of real data in a ciphertext and the matrix version of the $\operatorname{circshift}$ operation, which were introduced in the previous section.
@@ 4.4 Full algorithm for FHE-secured simulations
All introduced numerical schemes in Eqns. (16)–(19) are one-step methods. Alg. 3 represents the general structure of how all such numerical methods can be implemented with FHE. Here, it is assumed that the initial solution $\bm{u}^{0}$ is already encrypted. The algorithm iterates over the time span $t_{0}$ to $t_{\text{end}}$ with a fixed time step $\Delta t$. In each iteration, we check if during the next iteration (requiring $l_{\text{step}}$ multiplicative levels), the ciphertext level would would fall below one (or two, if iterative bootstrapping is used). If this is the case, the solution is bootstrapped to increase the ciphertext level again. The actual computation of the solution at the next time step is then performed in line 6. The algorithm is terminated when the final time $t_{\text{end}}$ is reached.
@@
1:$\bm{u}\leftarrow\bm{u^{0}}$triangleright Initialize solution
2:$t\leftarrow t_{0}$triangleright Initialize time
@@
4:$\bm{u}\leftarrow$bootstrap$(\bm{u})$EndIf
5:
6:$\bm{u}^{n+1}\leftarrow f(\bm{u})$triangleright Compute solution at next time step based on Eqns. (16)–(19)
7:$\bm{u}\leftarrow\bm{u}^{n+1}$
8:
[…]
@@
While the higher-order Lax-Wendroff scheme is more accurate in theory, it also uses more FHE operations per iteration, which increases both runtime and the error introduced by the CKKS scheme. Therefore, these requirements need to be balanced against each other in practice, and they likely depend on the specific application. A more detailed error and performance analysis for our particular setup will be presented in the next section.
@@ 5 Numerical results and performance of secure numerical simulations
For the numerical simulations presented here, the advection speed of the linear advection equations is set to $a_{x}=a_{y}=1$. The initial condition is given by $u_{0}(x)=\sin(2\pi x)$ in 1D and $u_{0}(x,y)=\sin(2\pi x)\sin(2\pi y)$ in 2D, and periodic boundary conditions are imposed. Unless noted otherwise, we will use $N=N_{x}=N_{y}=64$ equidistant nodes in each spatial direction to discretize the computational domain, and a simulation time span of $t\in[0,1]$. The time step size $\Delta t$ is determined from the CFL condition with a CFL number of $0.5$.
In the following, we will first conduct convergence tests to verify that the FHE-secured numerical simulations do not break the convergence properties of the numerical schemes, and present the numerical results of a full simulation. We will then discuss the accuracy and performance of the secure simulations and compare them to their unencrypted counterparts. Finally, we will briefly explore the potential for speeding up the simulations by leveraging the multi-threading capabilities of the OpenFHE library.
@@ 5.1 Convergence test
We begin by performing a convergence study to verify the implementation of the numerical methods and to make sure that the convergence properties of the numerical schemes are not affected by the FHE encryption. We will use the simulation setup as described, but vary the number of nodes in each direction from $N=32$ to $N=256$ and reduce simulation time span to $t\in[0,0.5]$. To speed up the convergence tests, we ran the simulations in parallel with eight threads (for more details on parallelizing OpenFHE with OpenMP, see Sec. 5.4 below).
@@
[…]
@@
1D, upwind $N$ $L^{2}$ error EOC 32 1.01e-01 - 64 5.25e-02 0.95 128 2.67e-02 0.97 256 1.35e-02 0.99 1D, Lax-Wendroff $N$ $L^{2}$ error EOC 32 1.07e-02 - 64 2.67e-03 2.00 128 6.69e-04 2.00 256 1.67e-04 2.00 2D, upwind $N$ $L^{2}$ error EOC 32 1.88e-01 - 64 1.07e-01 0.82 128 5.69e-02 0.90 256 2.94e-02 0.95 2D, Lax-Wendroff $N$ $L^{2}$ error EOC 32 1.07e-02 - 64 2.68e-03 2.00 128 6.69e-04 2.00 256 1.67e-04 2.00
Table 5: Convergence test results for the 1D and 2D discretizations of the linear scalar advection equation with the first-order upwind scheme and the second-order Lax-Wendroff scheme.
@@ 5.2 Results for a full simulation
With the convergence properties of the numerical solvers verified, we proceed with comparing the accuracy of the results and the performance of the secure algorithms with their unencrypted counterparts. We continue to use the same simulation setup as given above. Since all algorithms are implemented by using SecureArithmetic.jl, we can directly compare the results by running the simulations once with the OpenFHE/OpenFHE.jl backend and once with an unencrypted backend that internally uses plain Julia vectors/matrices. This allows us to measure the error and runtime behavior of the encrypted simulations relative to the unencrypted on otherwise identical implementations.
In Fig. 14 we analyze the accuracy of the cryptographically secure numerical simulations compared to the unencrypted simulations. Note that unlike the convergence test, where we used the exact solution as a reference, we now compare the results of the secure simulations to the results of the unencrypted simulations. This allows us to separate the error of the numerical discretization and instead only consider the error introduced by the FHE encryption, which we compute as the $L^{\infty}$ norm of the difference between the secure and unencrypted solutions at each time step. In both the 1D and 2D cases, we can see that the error for the Lax-Wendroff scheme is higher than for the upwind scheme. This is expected, since the Lax-Wendroff scheme requires more FHE operations, which in turn increases the error introduced by the CKKS scheme. As we saw above, however, overall the Lax-Wendroff scheme is still much more accurate than the upwind scheme, even with the additional error introduced by the encrypted operations. Furthermore, the relative $L^{\infty}$ error at $\mathcal{O}(10^{-6})$ is still several orders of magnitude lower then the absolute numerical error of the unencrypted simulations at $\mathcal{O}(10^{-3})$. Yet, it is also clear that with FHE, less than single-precision floating point accuracy is achieved, which may be a limiting factor for some applications.
@@
Next, we look at the runtime performance of the secure numerical simulations. In Table 6(a) we show the runtime per step for the different schemes, which in case of the secure simulations was measured between the first and the second bootstrapping operation, including bootstrapping, and then averaged over the number of time steps. From the table, we can clearly see that the secure simulations are significantly slower than the unencrypted simulations, by approximately six orders of magnitude. Considering the runtime measurements of the individual CKKS operations in Sec. 5.3, this is not surprising. Furthermore, the 2D simulations are more expensive than the 1D simulations, and the Lax-Wendroff scheme has a higher runtime than the upwind scheme, which also matches expectations. It is interesting to note that the runtime of the encrypted simulations only differs by a factor of two to four between the 1D and 2D simulations, while the unencrypted 2D simulations are around ten times more expensive than their 1D counterparts. Considering that the FHE operation count for the 2D simulations is much higher than for the 1D simulations, as presented in Table 4, this is a strong indication of the impact the bootstrapping operation has on the runtime of the secure simulations.
@@
To highlight their interdependence, in Fig. 16 we analyze the error and runtime of the secure simulations with the Lax-Wendroff scheme in 2D for different multiplicative depths after bootstrapping $l_{\text{refresh}}$. The error becomes smaller with larger $l_{\text{refresh}}$, aligning with our previous observation that the bootstrapping incurs a high approximation error. The overall runtime first decreases with increasing multiplicative depth, which is expected since the bootstrapping operation is less frequent. However, there seems to exist an optimum value, since after approximately $l_{\text{refresh}}=17$ the runtime increases again. This confirms our earlier discussion that there exists a trade-off between the computational cost of the FHE operations, which increases with larger values for $l_{\text{refresh}}$, and the frequency of the bootstrapping operation. The optimal multiplicative depth thus depends on the specific application and the used numerical scheme.
@@
Overall, our findings demonstrate that it is possible to implement secure numerical simulations with the CKKS scheme, and that we are able to obtain sufficiently accurate results with the implemented numerical methods for the linear advection equation. They also show that the CKKS scheme is a significant source of errors and that it introduces a considerable performance overhead. Especially the bootstrapping operation has a large impact on both error and runtime. For practical applications, it thus seems advisable to primarily focus on optimizing the multiplicative depth and the bootstrapping configuration to minimize the introduced error and to maximize the performance of secure simulations. In scenarios where runtime is not the primary limitation, iterative bootstrapping may be used to improve the accuracy of FHE-based simulations.
@@ 5.4 Parallel computations
Fig. 18 shows the runtimes for the parallel execution of the secure simulations. For the used CKKS parameters and hardware setup, we could achieve at most a three-fold speedup in computation time with eight threads, and a ten-fold speedup with $32$ threads in configuration time. In either case, using more threads than this did not lead to any significant improvements in runtime. Since the multithreading capabilities of the OpenFHE library are readily available and easy to use, it is advisable to employ them to speed up the CKKS computations whenever possible.
@@ 6 Extension to other mathematical models and numerical methods
As demonstrated in the previous section, secure numerical simulations are generally possible. It was also noted that the primary impediments to the practical use of the CKKS scheme for such scenarios are the large computational overhead and, to a lesser extent, the precision loss. Apart from this, three main challenges arise when considering to extend the simulation prototypes shown in this paper to other PDEs and to more complex numerical schemes: a single built-in data structure, limited function evaluation capabilities, and constraints in control flow logic when it comes to conditional branching. In the following, we will discuss these issues and potential solutions to them, to give other researchers a starting point for their own investigations into secure numerical simulations. We will conclude this section with a brief summary of some general strategies for improving performance.
@@ 6.1 Data structures
In general, the concrete choice of data representation depends on the target application and the used numerical method. Due to the significant computational overhead of the CKKS scheme, it is necessary to make full use of its batch processing capabilities, i.e., the ability to do element-wise addition and multiplication on the entire data vector. Therefore, discretization schemes that allow for efficient vectorization are preferable when considering other numerical methods for secure simulations. The choice of PDE is largely unaffected by the data structure, as long as a single model with a constant number of state variables is used throughout the computational domain.
@@ 6.2 Function evaluation
For many discontinuous functions, e.g., sign/comparison, the use of CKKS is challenging as the polynomial approximation in the proximity of the singularity point(s) requires extremely high-degree polynomials, i.e., very large multiplicative depth and runtime. Instead, other FHE schemes are often used, such as DM [23] or CGGI [24]. DM and CGGI support the evaluation of arbitrary functions via lookup tables using a technique called functional or programmable bootstrapping [59]. In these scenarios, a scheme switching procedure is employed [60, 61], where the encryption of the data is temporarily changed from CKKS to DM/CGGI. This allows for the evaluation of the discontinuous function, after which the data is converted back to a CKKS representation. Note that such scheme switching capability is supported in OpenFHE [40]. The main drawback of scheme switching is the significant computational overhead, especially when high precision is required [62, 51, 63]. A promising new method for evaluating discontinuous functions using functional CKKS bootstrapping is proposed and implemented in [64]. The method enables arbitrary function evaluation via lookup tables and achieves a throughput that is orders of magnitude higher than DM or CGGI.
6.3 Conditional branching
@@
As noted in [65], this restriction in control flow logic presents a fundamental challenge when converting regular programs to FHE algorithms. Since branching on an encrypted value is impossible by design, an FHE algorithm encountering a conditional branch must therefore first compute the outcome for each possible branch. The final result is then constructed by combining the results from all branches and applying an appropriate mask that selects the correct result. For example, Alg. 4 shows a simple branching statement in a regular program. Based on a function condition() that returns either $1$ or 00, the value of $r$ is computed by calling either foo() or bar(). The equivalent FHE algorithm in Alg. 5 first computes the value of the condition, evaluates both branches, and then combines the results with the condition value to obtain the final result.
@@
[…]
1:$r\leftarrow\text{foo}()$Else
2:$r\leftarrow\text{bar}()$EndIf
[…]
@@
1:$c\leftarrow\text{condition()}$()
@@
Such an approach is feasible if a small, predetermined number of branches needs to be considered, e.g., when computing the well-known HLLC flux for the compressible Euler equations, which has different branches based on three different wave speeds [66]. It becomes intractable, however, if the number of branches is large and possibly unknown in advance, e.g., when terminating a time step loop while using a dynamic, CFL-based time step size. Furthermore, numerical schemes that strictly require conditional branching based on current solution values are not directly compatible with FHE. This includes methods that dynamically modify their data structures during the simulation, such as periodic remeshing when dealing with deforming boundaries or adaptive mesh refinement. As a result, this poses a concrete limitation on the applicability of FHE to certain classes of numerical simulations.

### v3

@@ 4.3.3 Rewriting the finite difference methods in 2D
$\bm{u}^{n+1}$ is a matrix now, with coefficients $u_{ij}$ at each node location. Consequently, we also need to use the matrix representation of real data in a ciphertext and the matrix version of the $\operatorname{circshift}$ operation, which were introduced in the previous section.
@@ 4.4 Full algorithm for FHE-secured simulations
All introduced numerical schemes in Eqns. (16)–(19) are one-step methods. Alg. 3 represents the general structure of how all such numerical methods can be implemented with FHE. Here, it is assumed that the initial solution $\bm{u}^{0}$ is already encrypted. The algorithm iterates over the time span $t_{0}$ to $t_{\text{end}}$ with a fixed time step $\Delta t$. In each iteration, we check if during the next iteration (requiring $l_{\text{step}}$ multiplicative levels), the ciphertext level would would fall below one (or two, if iterative bootstrapping is used). If this is the case, the solution is bootstrapped to increase the ciphertext level again. The actual computation of the solution at the next time step is then performed in line 9. The algorithm is terminated when the final time $t_{\text{end}}$ is reached.
@@
1: $\bm{u}\leftarrow\bm{u^{0}}$ $\triangleright$ Initialize solution
2: $t\leftarrow t_{0}$ $\triangleright$ Initialize time 3:
4: while $t<t_{\text{end}}$ do
@@
6:    $\bm{u}\leftarrowbootstrap$$(\bm{u})$
7:   end if 8:
9:   $\bm{u}^{n+1}\leftarrow f(\bm{u})$ $\triangleright$ Compute solution at next time step based on Eqns. (16)–(19)
10:   $\bm{u}\leftarrow\bm{u}^{n+1}$ 11:
12:   $t\leftarrow t+\Delta t$ $\triangleright$ Update time
13: end while
@@
While the higher-order Lax-Wendroff scheme is more accurate in theory, it also uses more FHE operations per iteration, which increases both runtime and the error introduced by the CKKS scheme. Therefore, these requirements need to be balanced against each other in practice, and they likely depend on the specific application.
@@ 5 Numerical results and performance of secure numerical simulations
For the numerical simulations presented here, the advection speed of the linear advection equations is set to $a_{x}=a_{y}=1$. The initial condition is given by $u_{0}(x)=\sin(2\pi x)$ in 1D and $u_{0}(x,y)=\sin(2\pi x)\sin(2\pi y)$ in 2D, and periodic boundary conditions are imposed. Unless noted otherwise, we use $N=N_{x}=N_{y}=64$ equidistant nodes in each spatial direction to discretize the computational domain, and a simulation time span of $t\in[0,1]$. The time step size $\Delta t$ is determined from the CFL condition with a CFL number of $0.5$.
In the following, we first conduct convergence tests to verify that the FHE-secured numerical simulations do not break the convergence properties of the numerical schemes, and present the numerical results of a full simulation. We then discuss the accuracy and performance of the secure simulations and compare them to their unencrypted counterparts. Finally, we briefly explore the potential for speeding up the simulations by leveraging the multi-threading capabilities of the OpenFHE library.
@@ 5.1 Convergence test
We begin by performing a convergence study to verify the implementation of the numerical methods and to make sure that the convergence properties of the numerical schemes are not affected by the FHE encryption. We use the simulation setup as described, but vary the number of nodes in each direction from $N=32$ to $N=256$ and reduce the simulation time span to $t\in[0,0.5]$. To speed up the convergence tests, we ran the simulations in parallel with eight threads (for more details on parallelizing OpenFHE with OpenMP, see Sec. 5.4 below).
@@
$\text{1D: }\,e_{N}=\sqrt{\frac{1}{N}\sum_{i=1}^{N}(u_{i}^{n}-u_{\text{exact},i}^{n})^{2}}\qquad\text{2D: }\,e_{N}=\sqrt{\frac{1}{N^{2}}\sum_{i,j=1}^{N}(u_{ij}^{n}-u_{\text{exact},ij}^{n})^{2}}$ (20)
@@
1D, upwind NN $L^{2}$ error EOC 32 1.01e-01 - 64 5.25e-02 0.95 128 2.67e-02 0.97 256 1.35e-02 0.99 1D, Lax-Wendroff NN $L^{2}$ error EOC 32 1.07e-02 - 64 2.67e-03 2.00 128 6.69e-04 2.00 256 1.67e-04 2.00 2D, upwind NN $L^{2}$ error EOC 32 1.88e-01 - 64 1.07e-01 0.82 128 5.69e-02 0.90 256 2.94e-02 0.95 2D, Lax-Wendroff NN $L^{2}$ error EOC 32 1.07e-02 - 64 2.68e-03 2.00 128 6.69e-04 2.00 256 1.67e-04 2.00
Table 5: Convergence test results for the secure simulation of the 1D/2D linear scalar advection equation with the first-order upwind scheme and the second-order Lax-Wendroff scheme.
@@ 5.2 Results for a full simulation
With the convergence properties of the numerical solvers verified, we proceed with comparing the accuracy of the results and the performance of the secure algorithms with their unencrypted counterparts. In Fig. 16 we analyze the accuracy of the cryptographically secure numerical simulations compared to the unencrypted simulations, using the same setup as before. Unlike the convergence test, where we used the exact solution as a reference, we now compare the results of the secure simulations to the results of the unencrypted simulations. This allows us to separate the error of the numerical discretization and instead only consider the error introduced by the FHE encryption, which we compute as the $L^{\infty}$ norm of the difference between the secure and unencrypted solutions at each time step. In both the 1D and 2D cases, we can see that the error for the Lax-Wendroff scheme is higher than for the upwind scheme. This is expected, since the Lax-Wendroff scheme requires more FHE operations, which in turn increases the error introduced by the CKKS scheme. As we saw above, however, overall the Lax-Wendroff scheme is still much more accurate than the upwind scheme, even with the additional error introduced by the encrypted operations. Furthermore, the relative $L^{\infty}$ error at $\mathcal{O}(10^{-6})$ is still several orders of magnitude lower then the absolute numerical error of the unencrypted simulations at $\mathcal{O}(10^{-3})$. Yet, it is also clear that with FHE, less than single-precision floating point accuracy is achieved, which may be a limiting factor for some applications.
@@
Next, we look at the runtime performance of the secure numerical simulations. In Table 6(a) we show the runtime per step for the different schemes, which in case of the secure simulations was measured between the first and the second bootstrapping operation, including bootstrapping, and then averaged over the number of time steps. From the table, we can clearly see that the secure simulations are significantly slower than the unencrypted simulations, by approximately six orders of magnitude. Considering the runtime measurements of the individual CKKS operations in Sec. 5.3, this is not surprising. Furthermore, the 2D simulations are more expensive than the 1D simulations, and the Lax-Wendroff scheme has a higher runtime than the upwind scheme, which also matches expectations. It is interesting to note that the runtime of the encrypted simulations only differs by a factor of two to four between the 1D and 2D simulations, while the unencrypted 2D simulations are around ten times more expensive than their 1D counterparts. Given that the FHE operation count in the 2D simulations is considerably higher than in 1D (Table 4), the relatively modest increase in runtime suggests that bootstrapping dominates the overall performance cost. This is further supported by the measured bootstrapping time: for the Lax-Wendroff scheme, bootstrapping accounts for about 54% of the total runtime in 1D and 36% in 2D.
@@
To highlight their interdependence, in Fig. 18 we analyze the error and runtime of the secure simulations with the Lax-Wendroff scheme in 2D for different multiplicative depths after bootstrapping $l_{\text{refresh}}$. The error becomes smaller with larger $l_{\text{refresh}}$, aligning with our previous observation that the bootstrapping incurs a high approximation error. The overall runtime first decreases with increasing multiplicative depth, which is expected since the bootstrapping operation occurs less frequently. However, there seems to exist an optimum value, since after approximately $l_{\text{refresh}}=17$ the runtime increases again. This confirms our earlier discussion that there exists a trade-off between the computational cost of the FHE operations, which increases with larger values for $l_{\text{refresh}}$, and the frequency of the bootstrapping operation. The optimal multiplicative depth thus depends on the specific application and the used numerical scheme.
@@
Overall, our findings demonstrate that secure numerical simulations with the CKKS scheme are feasible. Results remain accurate as long as bootstrapping is not required, which introduces noticeable numerical errors—though single-precision accuracy is still achievable. The results also highlight the considerable performance overhead introduced by CKKS—primarily from bootstrapping. For practical applications, it thus seems advisable to primarily focus on optimizing the multiplicative depth and the bootstrapping configuration to minimize the introduced error and to maximize the computational performance of secure simulations. In scenarios where runtime is not the primary limitation, iterative bootstrapping may be used to improve the accuracy of FHE-based simulations.
@@ 5.4 Parallel computations
Fig. 20 shows the runtimes for the parallel execution of the secure simulations. For the used CKKS parameters and hardware setup, we could achieve at most a three-fold speedup in computation time with eight threads, and a ten-fold speedup with $32$ threads in configuration time. In either case, using more threads than this did not lead to any significant improvements in runtime.
Given that using multiple threads can reduce solution times and that the OpenFHE library offers basic multithreading capabilities, it is advisable to employ 2 to 8 threads to speed up CKKS computations whenever possible. However, the observed speedup remains fall short of ideal scaling, indicating that the current multithreading implementation has practical limitations. Since efficient parallelization is crucial for the viability of FHE-based simulations—especially in more demanding applications—this remains an important subject for future investigation. Nevertheless, these results provide a basis for exploring more advanced simulation scenarios, outlined in the next section.
@@ 6 Extension to other mathematical models and numerical methods
Building on the prototype simulations and findings from Sec. 5, we discuss challenges and possible strategies for extending secure numerical simulations to more complex models and schemes. As noted before, the primary impediments to the practical use of the CKKS scheme for such scenarios are the large computational overhead and, to a lesser extent, the precision loss. In addition, three main challenges arise when considering to extend the simulation prototypes shown in this paper to other PDEs and advanced numerical schemes: a single built-in data structure, limited function evaluation capabilities, and constraints in control flow logic when it comes to conditional branching. In the following, we discuss these issues and potential solutions to them, to give other researchers a starting point for their own investigations into secure numerical simulations. We then outline general strategies to improve performance and conclude with a brief summary of simulation types that are (currently) out of scope.
@@ 6.1 Data structures
In general, the concrete choice of data representation depends on the target application and the used numerical method. Due to the significant computational overhead of the CKKS scheme, it is necessary to make full use of its batch processing capabilities, i.e., the ability to do element-wise addition and multiplication on the entire data vector. Therefore, discretization schemes that allow for efficient vectorization are preferable when considering other numerical methods for secure simulations. The choice of PDE is largely unaffected by the data structure, as long as a single model with a constant number of state variables is used throughout the computational domain. However, algorithms that require fine-grained access to the data—such as irregular sparse matrix operations—are not directly compatible with FHE.
@@ 6.2 Function evaluation
For many discontinuous functions, e.g., sign/comparison, the use of CKKS is challenging as the polynomial approximation in the proximity of the singularity point(s) requires extremely high-degree polynomials, i.e., very large multiplicative depth and runtime. Instead, other FHE schemes are often used, such as DM [23] or CGGI [24]. DM and CGGI support the evaluation of arbitrary functions via lookup tables using a technique called functional or programmable bootstrapping [61]. In these scenarios, a scheme switching procedure is employed [62, 63], where the encryption of the data is temporarily changed from CKKS to DM/CGGI. This allows for the evaluation of the discontinuous function, after which the data is converted back to a CKKS representation. Such scheme switching capability is supported in OpenFHE [40]. The main drawback of scheme switching is the significant computational overhead, especially when high precision is required [64, 52, 65]. A promising new method for evaluating discontinuous functions using functional CKKS bootstrapping is proposed and implemented in [53]. The method enables arbitrary function evaluation via lookup tables and achieves a throughput that is orders of magnitude higher than DM or CGGI.
6.3 Data-dependent control flow
@@
As noted in [66], this restriction in control flow logic presents a fundamental challenge when converting regular programs to FHE algorithms. Since branching on an encrypted value is impossible by design, an FHE algorithm encountering a conditional branch must therefore first compute the outcome for each possible branch. The final result is then constructed by combining the results from all branches and applying an appropriate mask that selects the correct result. For example, Alg. 4 shows a simple branching statement in a regular program. Based on a function condition() that returns either $1$ or 0, the value of $r$ is computed by calling either foo() or bar(). The equivalent FHE algorithm in Alg. 5 first computes the value of the condition, evaluates both branches, and then combines the results with the condition value to obtain the final result.
@@
1: if $\text{condition}()==1$ then
2:   $r\leftarrow\text{foo}()$
3: else
4:   $r\leftarrow\text{bar}()$
5: end if
@@
1: $c\leftarrow\text{condition()}$
@@
Such an approach is feasible if a small, predetermined number of branches needs to be considered, e.g., when computing the well-known HLLC flux for the compressible Euler equations, which has different branches based on three different wave speeds [67]. It becomes intractable, however, if the number of branches is large and possibly unknown in advance, e.g., when terminating a time step loop while using a dynamic, CFL-based time step size or when considering iterative schemes with error-based stopping criteria. Furthermore, numerical schemes that strictly require conditional branching based on current solution values are not directly compatible with FHE. This includes methods that dynamically modify their data structures during the simulation, such as periodic remeshing when dealing with deforming boundaries or adaptive mesh refinement. As a result, this poses a concrete limitation on the applicability of FHE to certain classes of numerical simulations.